OpenStudy (anonymous):

Solve the system of equations: 9x-3y=3 3x+8y=-17

OpenStudy (anonymous):

you can solve this system using many methods. multiply equation 2 by -3 and add the 2 equations to eliminate x

OpenStudy (anonymous):

9x-3y=3 -9x-24y=51 =>-27y=54=>y=-2 now substitute y=-2 in one of the equations: 9x+6=3 9x=-3 =>x=-1/3 (x;y)=(-1/3;-2)

OpenStudy (anonymous):

so x is -1/3 and y is -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks